OpenStudy (anonymous):

If someone can, I need someone to solve this so I can compare my answer with their's on the quadratic formula.

OpenStudy (anonymous):

\[2x^2-x-2=0\] My answer was: \[x+/-1i \sqrt{x^2}/2\]

OpenStudy (anonymous):

If you can't read that end piece, it's radical x^2 over 2

OpenStudy (accessdenied):

how did you get that answer? quadratic formula doesn't involve x's in it

OpenStudy (anonymous):

Heh... I miraculously just caught that. Give me a bit.

OpenStudy (anonymous):

For the b in this equation, it would be -1 correct?

OpenStudy (accessdenied):

yes

OpenStudy (radar):

and a=2, c=-2 and you correctly identified the value of b

OpenStudy (anonymous):

So when I substituted b^2-4ac with \[-1^2-4(2)(-2)\], I should get \[\sqrt{1+16}\] Correct?

OpenStudy (anonymous):

As well as the 2(2) being 4 on the denominator.

OpenStudy (accessdenied):

yep. (tho i recommend using parentheses around the -1 when squaring it, since the exponent applies to the 1 in -1)

OpenStudy (anonymous):

I cannot simply this equation any further can I?

OpenStudy (radar):

\[1\pm \sqrt{1+16}\over4\]

OpenStudy (anonymous):

Since the 17 in the radical can't be rooted, that should be the solution?

OpenStudy (anonymous):

Unless that 4 divides into the 16.

OpenStudy (accessdenied):

yes, you cannot simplify it any further

OpenStudy (radar):

\[1\pm \sqrt{17}\over 4\]

OpenStudy (radar):

Don't forget the "-b" in the very beginning of the numerator.

OpenStudy (anonymous):

Thank you kindly.

OpenStudy (radar):

@jpigott, did you get my post regarding the -b in the numerator??

OpenStudy (anonymous):

Well i'm not necessarily sure what is supposed to happen to it. If I were to take a crack at it would it be 1+17/4 and 1-17/4?

OpenStudy (radar):

The final answer could look like:\[1\pm \sqrt{17}\over 4\]

OpenStudy (radar):

or 1/4 +/- (1/4)sqrt 17

OpenStudy (radar):

I just didn't want you to just drop it, it is part of the solution.

OpenStudy (radar):

Your last post is partially correct, but the 17 should be sqrt 17, or shown in a radical.

OpenStudy (radar):

The equation editor is handy for expressing quadratics and their solution. All you have to do is click on the left lower button below your typing box.

OpenStudy (radar):

The one labeled "Equation"\[\Sigma"Equation"\]

OpenStudy (anonymous):

I do that. My apologies for the confusion. If i am reading what you typed correctly, I have to do (1+√17)/4 and vise vera?

OpenStudy (radar):

By Jove, I believe you have it. Good luck with them.

OpenStudy (anonymous):

Oh now I see what you were trying to get at. The actual answer is supposed to include both the + and the 1 equation as the answer, right?

OpenStudy (radar):

Yes,\[1+\sqrt{17}\over 4\] and \[1-\sqrt{17}\over 4\]

OpenStudy (anonymous):

Whoop! Thank you!

OpenStudy (radar):

U r welcome.